# Machine Learning and the Physical Sciences

### v1

@@ Appendix B Tailwise conformal prediction intervals
In the more basic forms of conformal prediction, uncertainty intervals are constructed symmetrically—predictive errors above and below are assumed to have the same distribution. Concretely, for a model $f:X\rightarrow Y$, if we denote the residual
@@
The aforementioned construction is simple and guarantees the desired coverage over the calibration data, but it forces the lower and upper bounds to be equally wide. Hence, following the tailwise quantile approach first introduced in [barber_predictive_2021], we use tailwise (one-sided) quantiles. For a desired miscoverage $\alpha$, we split the miscoverage evenly between the tails on both sides, with $\alpha/2$ mass for each. We then compute the lower and upper quantiles for the signed residuals: $Q_{\alpha/2}(R)$ and $Q_{1-\alpha/2}(R)$. This defines the prediction intervals used in this study:
@@
which still guarantees $1-\alpha$ coverage over the validation set but allows the lower and upper margins to differ whenever the residual distribution is asymmetric.
@@ Appendix C Reproducibility Details for AE-SINDy Model Training
This appendix provides the essential information required to reproduce the autoencoder-SINDy (AE-SINDy) model setup and training procedure used in this study. The description covers data acquisition and preprocessing, model architecture and hyperparameters, and the training workflow, including loss function specification. The source code and data are further included in a linked repository.
@@ C.1 Data Source and Preprocessing
The AE-SINDy model is trained using binned particle size distribution (PSD) data generated from a large-eddy simulation (LES) employing the superdroplet method. We simulate the evolution of a warm liquid-phase cloud that forms from a gaussian surface moisture and heat flux, growing in altitude before precipitating. The primary datasets are accessed in NetCDF format and contain variables for binned droplet mass distributions over time and space with droplet coalescence active as the only enable droplet dynamic. For model input, only samples with sufficient liquid water content (e.g., exceeding $10^{-5}$ kg/kg) are included. Each PSD is normalized by its total liquid mass to ensure scale invariance during encoding.
@@ C.2 Model Architecture and Hyperparameters
The AE-SINDy model consists of three main components:
@@
    Encoder: A feed-forward neural network (FFNN) with four fully-connected layers, sequentially reducing the input dimension from $N_{\text{bins}}$ to the latent dimension (minus the total mass variable), i.e. $N_{\text{latent}}-1=3$. ReLU activations are used on hidden layers, with the final layer mapping to the latent space without nonlinearity.
@@
    Decoder: A FFNN mirroring the encoder structure, with four fully-connected layers expanding from $N_{\text{latent}}-1$ back to $N_{\text{bins}}$. The output layer applies a softmax activation to ensure the reconstructed PSD is normalized.
@@
    SINDy Dynamics Module: A bias-free, single-layer neural network representing the sparse identification of nonlinear dynamics (SINDy). This layer outputs a linear combination of time derivatives in the latent space using a polynomial feature library up to order 2.
@@ C.3 Training Procedure and Loss Function
Training is performed using the AdamW optimizer with learning rate scheduling and early stopping based on validation loss. The model is trained for up to 1000 epochs, with the option to halt training if no improvement is observed over a specified patience interval.
@@
Loss weights are chosen to balance the reconstruction and dynamics learning, following a scaling based on the relative magnitudes of the PSD and its time derivative in the training data as in [champion_data-driven_2019]. Other parameters, including the batch size, initial learning rate, and a multiplicative factor of $w_{dx}$, were determined using hyperparameter optimization.
@@ C.4 Code Availability
All code used for data processing, model definition, and training is written in Python using PyTorch and is available at https://anonymous.4open.science/r/UQ_AE-SINDy-1A36. The scripts include utilities for loading NetCDF datasets, constructing PyTorch DataLoaders, defining the AE-SINDy architecture, executing the training loop, and running and visualizing the uncertainty quantification pipelines.

### v2

@@ Appendix B Nonconformity scores
In the more basic forms of conformal prediction, uncertainty intervals are constructed symmetrically—predictive errors above and below are assumed to have the same distribution. Concretely, for a model $f:X\rightarrow Y$, if we denote the signed residual
@@
The aforementioned construction is simple and guarantees the desired coverage over the calibration data, but it forces the lower and upper bounds to be equally wide. Hence, following the tailwise quantile approach first introduced in [barber_predictive_2021], we use tailwise (one-sided) quantiles for the DSD-valued predictions. For a desired miscoverage $\alpha$, we split the miscoverage evenly between the tails on both sides, with $\alpha/2$ mass for each. We then compute the lower and upper quantiles for the signed residuals: $Q_{\alpha/2}(R)$ and $Q_{1-\alpha/2}(R)$. This defines the prediction intervals used in this study:
@@
which still ensures $1-\alpha$ coverage over the validation set but allows the lower and upper margins to differ whenever the residual distribution is asymmetric.
For latent-space predictions, the outputs are multivariate and typically exhibit correlations across dimensions. Hence, in this case, coordinate-wise absolute or signed residuals may not accurately capture the joint error structure. Accordingly, for a predicted latent vector $\hat{z}=f(x)$ and true latent vector $z$, we define the residual $r=z-\hat{z}$ and use a scalar nonconformity score based on the (squared) Mahalanobis distance:
$S(z,\hat{z})=r^{\top}\Sigma_{r}^{-1}r,$
where $\Sigma_{r}$ is the empirical covariance matrix of latent-space residuals estimated from the calibration set (in practice, using a Ledoit–Wolf shrinkage estimator [LedoitWolf2012, scikit-learn]). In particular, the scalar score S(z,z^)S($\hat{z}$) is used to compute the empirical $(1-\alpha)$-quantile on the calibration set, which then defines the size of the latent-space prediction ellipsoid. Using the residual covariance rather than the latent covariance yields a conformity score that is properly normalized with respect to the model’s error geometry and captures correlated uncertainty through multivariate prediction ellipsoids.
The prediction intervals and nonconformity scores described above are computed independently at each timestep, using instantaneous residuals and covariance estimates. Hence, the prediction sets evolve in time, therby reflecting the temporal evolution of model uncertainty.
@@ Appendix C Reproducibility Details for AE-SINDy Model Training
This appendix provides the essential information required to reproduce the autoencoder-SINDy (AE-SINDy) model setup and training procedure used in this study. The description covers data acquisition and preprocessing, model architecture and hyperparameters, and the training workflow, including loss function specification. The source code and data are further included in a linked repository. Further details on the AE-SINDy architecture, the polynomial SINDy library, and the loss structure used during training are provided in the companion study De Jong et al. [dejong2025data], which develops the surrogate model in full.
@@ C.1 Data Source and Preprocessing
The AE-SINDy model is trained using binned particle size distribution (PSD) data generated from a large-eddy simulation (LES) employing the superdroplet method. We simulate the evolution of a warm liquid-phase cloud that forms from a Gaussian surface moisture and heat flux, growing in altitude before precipitating. The primary datasets are accessed in NetCDF format and contain variables for binned droplet mass distributions over time and space, with droplet coalescence active as the only enabled droplet dynamic. For model input, only samples with sufficient liquid water content (e.g., exceeding $10^{-5}$ kg/kg) are included. Each PSD is normalized by its total liquid mass to ensure scale invariance during encoding.
@@
For a full description of the large eddy simulations used to generate the PSD training data—including the simulation setups, temporal sampling, grid resolution, and the physics assumed in the forward simulation—see De Jong et al. [dejong2025data], which details the simulation design and post-processing used to produce the binned PSD datasets.
@@ C.2 Model Architecture and Hyperparameters
The AE-SINDy architecture used here follows the structure introduced in [dejong2025data], which employs four fully-connected encoder and decoder layers that halve (or double, respectively) the dimension at each layer, using ReLU activations in the hidden layers and a softmax output to ensure normalized reconstructions, and employs a SINDy model for the dynamics in the latent space. More explicitly, the model consists of three components:
@@
    Encoder: A feed-forward neural network (FFNN) with four fully connected layers that sequentially reduce the input dimension from $N_{\text{bins}}$ to the latent dimension (excluding the total-mass variable), yielding $N_{\text{latent}}-1=3$ latent variables. Hidden layers use ReLU activations, and the final layer maps to the latent space without a nonlinearity.
@@
    Decoder: A FFNN mirroring the encoder structure, with four fully connected layers expanding from $N_{\text{latent}}-1$ back to $N_{\text{bins}}$. A softmax activation on the output layer enforces that reconstructed PSDs remain normalized.
@@
    SINDy Dynamics Module: A bias-free, single-layer neural network that implements the SINDy formulation by outputting a linear combination of latent-space time derivatives using a polynomial feature library of terms up to second order.
@@ C.3 Training Procedure and Loss Function
Training procedures for the AE-SINDy architecture follow the approach detailed in [dejong2025data] and the associated code repository, including early stopping, learning-rate scheduling, and a composite loss combining reconstruction, PSD-derivative, and latent-derivative terms.
In summary, training is performed using the AdamW optimizer with learning rate scheduling and early stopping based on validation loss. The model is trained for up to 1000 epochs, with the option to halt training if no improvement is observed over a specified patience interval.
@@
Loss weights are chosen to balance the reconstruction and dynamics learning, following a scaling based on the relative magnitudes of the PSD and its time derivative in the training data as in Champion et al. [champion_data-driven_2019]. Other parameters, including the batch size, initial learning rate, and a multiplicative factor of $w_{dx}$, were determined using hyperparameter optimization with Optuna [Optuna].
@@ C.4 Code Availability
All code used for data processing, model definition, and training is written in Python using PyTorch and is available at https://github.com/jonaskat87/UQ_AE-SINDy. The scripts include utilities for loading NetCDF datasets, constructing PyTorch DataLoaders, defining the AE-SINDy architecture, executing the training loop, and running and visualizing the uncertainty quantification pipelines.
Additional scripts for data processing, model specification, and training of the AE-SINDy surrogate are available in the code repository accompanying the companion paper De Jong et al. [dejong2025data].